OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (1, 5) and (-2, 3) The line that passes through (1, 5) and (-2, 3)

OpenStudy (anonymous):

To find the slope you can do change in y over change in x which would be 2/3 which is the slope. So now the equation becomes y = 2/3x + b to find be plug any of the points given in to the equation lets pick the first one for convenience 5 = 2/3*1 + b Now algebra... 15/3 - 2/3 = 13/3 = b I don't know how your teacher likes fractions so if you wanted it in mixed number it would be 4 1/3 So now convert y = 2/3x + 13/3 into standard form. -2/3x + y = 13/3 now multiply by -3 and you get 2x - 3y = -13 and that's your answer